OpenStudy (asrd98):

what is the solution to the inequality |2n+5|>1? -3>n>-2 2-2 n<2 or n>3

OpenStudy (sooobored):

\[|x|>c\] is the equivalent of saying x>c and -x>c

OpenStudy (amorfide):

|x|=x |-x|=x means, whether it is f(x) or -f(x), |f(x)|=|-f(x)| so you have a function in n, f(n)=2n+5 you want to know when it is greater than 1, so if we let f(n)=1 we get 2n+5=1 also, we know that we can let -f(n)=1, since |f(n)|=|-f(n)|=1 so -(2n+5)=1 solve both for n 2n+5=1, n=-4/2=-2 -(2n+5)=1 -2n-5=1 -6=2n -6/2=n n=-3 so we know that n=-2, or n=-3 now we want to know what values of n, |2n+5|>1 we know that when n=-2, |2n+5|=1, and when n=-3, |2n+5|=1, now we would usually use a graph to show you this  from the shaded region, we can easily see that when n is between -2 and -3, that is, -3=<n<=-2, we have |f(n)|<=1, so we want |f(n)|>1, so we want where the function is above y=1 therefore, your answer will be n>-2, or n<-3